Using Algebra Suppose M is between L and N. Use the Segment Addition Postulate to solve the variable. Then find the lengths of LM MN and LN
LM=3x+8
MN=2x-5
LN=23

Answer by checkley75(3666)

LM+MN=23 THEN
3X+8+2X-5=23
5X+3=23
5X=20
X=20/5
X=4 THEREFORE:
LM=3*4+8
LM=12+8
LM=20 ANSWER.
MN=2*4-5
MN=8-5
MN=3 ANSWER.